If the demand curve for comic books is expressed as Q=50,000/p, then compute the price elasticity of demand.


1
Expert's answer
2020-10-21T11:24:43-0400

we have "Q = \dfrac{50000}{p}"

As elasticity of demand "=\dfrac{dQ}{dP}*\dfrac{P}{Q}"

As "\dfrac{dQ}{dP} =\dfrac{d}{dP}(\dfrac{50000}{P})"

"=\dfrac{-50000}{P^2}"

"\therefore" price elasticity is "=\dfrac{-50000}{P^2}* \dfrac{P}{Q}"

so the right answer is "=\dfrac{-50000}{PQ}"

AS PQ is given as 50000

"\therefore" elasticity "==\dfrac{-50000}{50000} = -1"

hence the elasticity is unit unitary